a walkway is 10 feet wide which is going around a pond that is 200 feet wide. what will be the area, in square fee of the walkway.
Answer by richwmiller(9144)

find the area of a circle 220 wide. ie radius 110
find the area of a circle 200 wide ir radius 100
a=3.14*110^2
b=3.14*100^2
a = 37994 b = 31400
a-b
subtract a-b
a-b=6594